Country A has 100 workers and Country B has 100 workers. Every worker in Country A can produce 6 tons of wheat per year, or can produce 12 tons of corn per year. Every worker in Country B can produce 2 tons of wheat per year, or can produce 10 tons of corn per year.

Suppose initially the countries do not trade and Country A has 50 workers producing corn and 50 producing wheat. Country B has 30 workers producing corn and 70 producing wheat. Fill out the following table for Country B & PLEASE SHOW ME HOW YOU FOUND THE ANSWER FOR COUNTRY B.

Country A Country B
Corn Produced 600 ?
Wheat Produced 300 ?
Answer #1

In case of country A,

Wheat produced=output of wheat per worker*workers employed=6*50=300 tons

Corn produced=output of corn per worker*workers employed=12*50=600 tons

In case of country B,

Wheat produced=output of wheat per worker*workers employed=2*70=140 tons

Corn produced=output of corn per worker*workers employed=10*30=300 tons

Country A Country B
Corn Produced 600 300
Wheat Produced 300 140